OpenStudy (anonymous):

somebody please help! what is the equation of the line in standard form that passes through the point (-1,4) and is parallel to the line y=2x-3

OpenStudy (allank):

the gradients of parallel lines are equal, thus the gradient of the unknown line is 2 the equation of the line in point-slope form is y=2(x-(-1))+4 y=2(x+1)+4 y=2x+6 rewrite in standard form: 2x-y=-6

OpenStudy (anonymous):

ok ,, :) thanks

OpenStudy (allank):

you're welcome :)